OpenStudy (anonymous):

I need help with understanding what the y is in fractions and if you can show me how to get this answer 3y = 15x-9

OpenStudy (anonymous):

you want to know what the y intercept is ? or do you want put this equation in y= mx+b format?

OpenStudy (anonymous):

I want to know what the y intercept is. please.

OpenStudy (amistre64):

divide it all by 3 and make x=0

OpenStudy (amistre64):

 the y axis IS the x=0 axis

OpenStudy (amistre64):

so the yintercept is also the x=0 intercept

OpenStudy (anonymous):

or you can just set x to be zero like this: 3y=15(0)-9, then you get: 3y=-9, finally divide both sides by 3 and get : y=-3, thus y intercept is (0,-3)

OpenStudy (anonymous):

can you help me with this to please and thanks for the other answers and help 6y-3x=12

OpenStudy (anonymous):

to find the y intercept, simply set x to be zero like so: 6y-3(0)=12, now evluate that

OpenStudy (anonymous):